OpenStudy (anonymous):

Will someone please explain to me the formula for a car loan 18,708 With a payment of 406.76 per month at 11% intrest per year is 60 Months?

OpenStudy (anonymous):

will someone please help?

OpenStudy (anonymous):

Will someone please explain to me the formula for a car loan 18,708 With a payment of 406.76 per month at 11% intrest per year is 60 Months?

OpenStudy (anonymous):

Is the interested payed out monthly?

OpenStudy (anonymous):

As in, do you get 1/12 of the interest rate compounded every month?

OpenStudy (anonymous):

I am unsure. the 60 months is the only thing thats not in the previous statement... and the part asking for someone to explain... sorry, i don't know

OpenStudy (dumbcow):

formula involves Present Value monthly interest = .11/12 present value (v) of 1 dollar after 1 month = 1/(1+.11/12) = .9909166 \[18,708 = 406.76(v+v^{2}+...v^{60})\] using sum of geometric series \[18,708 = 406.76\frac{v(1-v^{60})}{1-v}\]

OpenStudy (shane_b):

I believe the formula you're looking for is: \[\large \frac{P(\frac{r}{12})}{1-(1+\frac{r}{12})^{-m}}\] \[\large \frac{$18708(\frac{0.11}{12})}{1-(1+\frac{0.11}{12})^{-60}}=$406.76\]